OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. 3x - 7y - z = -14 3x + 3y + 2z = 1 2x + 5y - 2z = -16 a. 3x - 7y = -14 2x + 5y = -16 b. 9x - 11y = -27 5x + 8y = -15 c. 9x - 11y = -27 3x - 7y = -14 d. 2x + 5y = -16 5x + 8y = -15

OpenStudy (anonymous):

B yet again 6x - 14y -2 z = -28 3x + 3y + 2z = 1 3x + 3y + 2z = 1 2x + 5y - 2z = -16 9x-11y=-27 5x+8y=-15 B B B

OpenStudy (anonymous):

Oh, crap. Sorry -- I asked the wrong one >> Ugh.

OpenStudy (anonymous):

Then my friend what is the right one? It is often said that we have to ask the wrong questions to get the right answers.